# Efficient Modelling of Lyman-$\alpha$ opacity fluctuations during late EoR

### v1

@@
Efficient Modelling of Lyman-$\alpha$ opacity fluctuations during late EoR
Barun Maity, maity@mpia.de11    Frederick Davies 11       Prakash Gaikwad 22
@@
The Lyman-$\alpha$ forest opacity fluctuations observed from high redshift quasar spectra have been proven to be extremely successful in order to probe the late phase of the reionization epoch. For ideal modeling of these opacity fluctuations, one of the main challenges is to satisfy the extremely high dynamic range requirements of the simulation box, resolving the Lyman alpha forest while probing the large cosmological scales. In this study, we adopt an efficient approach to model Lyman-$\alpha$ opacity fluctuations in coarse simulation volume, utilizing the semi-numerical reionization model SCRIPT (including inhomogeneous recombination and radiative feedback) integrated with a realistic photoionization background fluctuation generating model. Our model crucially incorporates ionization and temperature fluctuations, which are consistent with the reionization model. After calibrating our method with respect to high-resolution full hydrodynamic simulation, Nyx, we compare the models with available observational data at the redshift range, $z=5.0-6.1$. With a fiducial reionization model (reionization end at $z=5.8$), we demonstrate that the observed scatter in the effective optical depth can be matched reasonably well by tuning the free parameters of our model i.e. the effective ionizing photon mean free path and mean photoionization rate. We further pursue an MCMC-based parameter space exploration utilizing the available data to put constraints on the above free parameters. Our estimation prefers a slightly higher photoionization rate and slightly lower mean free path than the previous studies which is also a consequence of temperature fluctuations. This study holds significant promise for efficiently extracting important physical information about the Epoch of Reionization, utilizing the wealth of available and upcoming observational data.
@@ 1 Introduction
The Epoch of Reionization (EoR) represents an important phase in the history of the Universe, marking the time when the first stars and galaxies form and ionize neutral hydrogen atoms (HI) in the intergalactic medium (IGM) (see reviews for details, Barkana & Loeb, 2001; Choudhury, 2009; McQuinn, 2016; Dayal & Ferrara, 2018; Gnedin & Madau, 2022; Choudhury, 2022). The precise timeline and morphology of this epoch can be crucial to understand the structure formation and evolution of the universe at those early times. There exist a handful of observational probes at various wavelengths to trace the IGM during the EoR (Zaroubi, 2013; Mesinger, 2016). Among these, spectra obtained from high redshift quasars have been proven to be a treasure trove in extracting critical information about the nature of the IGM.
@@
Specifically, Lyman-$\alpha$ (Ly-$\alpha$) forest absorption of quasar spectra contains information about the line of sight distribution of neutral hydrogen atoms (HI) in the IGM, which has been widely used in the literature to study various physical aspects. For example, these include exploiting realistic numerical simulations to constrain the photoionization rate (Bolton & Haehnelt, 2007; Wyithe & Bolton, 2011; Gaikwad et al., 2023; Davies et al., 2024b) and tracking the temperature evolution of the IGM (Walther et al., 2019; Gaikwad et al., 2020). The high quality of existing quasar spectra also allows us to directly estimate the neutral fraction of the universe with damping wing analysis (Spina et al., 2024) and with dark pixel analysis (McGreer et al., 2015; Jin et al., 2023). Recently, a consensus of late reionization end ($z_{\mathrm{end}}\leq 5.3$) has come out based on the findings of large Gunn-Peterson (GP) troughs, giving rise to excess Ly-$\alpha$ forest opacity fluctuations at $z\geq 5.3$ (Zhu et al., 2021, 2022; Bosman et al., 2022).
@@
Several traditional models have been found to be inconsistent in order to explain the large GP trough and the excess Ly-$\alpha$ opacity fluctuations. There exist a few suggestions to match the data keeping the reionization end at $z\gtrsim 6$. One of these includes a hidden abundance of faint quasar populations at $z\geq 5$ (Meiksin, 2020), however, it is not yet completely understood whether they have the required photon budget to make a significant contribution. An extended and hot reionization scenario can also produce significant fluctuations via the temperature dependence of recombination rate (D’Aloisio et al., 2015). But, this model associates the opaque trough with large scale galaxy overdensity (Davies et al., 2018a) which is disfavoured by the galaxy survey results (Becker et al., 2018; Christenson et al., 2021). It is also possible that the excess ionizing background fluctuations are amplified due to very short mean free path of the ionizing photons in the IGM (Davies & Furlanetto, 2016; D’Aloisio et al., 2018). However, that doesn’t agree with the extrapolation from existing lower redshift measurements (Worseck et al., 2014). Another emerging explanation involves a late reionization end ($z_{\mathrm{end}}\leq 5.5$) where the shadowing effect on the large scale ionization fluctuations by the residual neutral islands is responsible for the excess opacity fluctuations (Kulkarni et al., 2019; Keating et al., 2020; Nasir & D’Aloisio, 2020). However, these models are finely tuned to match the ionizing photon emissivity according to the demand of both Ly-$\alpha$ forest fluctuations and other high redshift observables like CMB scattering optical depth. In recent times, this explanation has also been adopted by semi numerical models which further able to provide constraints on the global ionization fraction utilizing Ly-$\alpha$ forest observations (Choudhury et al., 2021; Qin et al., 2021, 2024). However, these models lack a realistic description of photoionization background fluctuations and temperature fluctuations. Despite recent improvements along these lines (Choudhury & Chakraborty, 2025), those models are still devoid of proper calibration against a realistic forest model.
@@
Nevertheless, it is crucial to determine the strength of photoionization background fluctuations as it provides an important boundary limit to the reionization scenario. It has been shown that the fluctuations rise rapidly when the ionization bubbles start to overlap with each other eliminating the neutral hydrogen patches (Bolton & Haehnelt, 2007; Wyithe & Bolton, 2011; Davies et al., 2018d). The measurements of the photoionization rate ($\Gamma_{\mathrm{HI}}$) also show an increment in the redshift range $z\sim 5-6$ supporting the above statement. However, the constraints are not very stringent due to lack of transmission flux at these high redshifts and a degeneracy with the thermal state of the IGM. In recent studies, the constraints have become more stringent with the abundance of high quality data and robust modelling (Gaikwad et al., 2023; Davies et al., 2024b).
@@
Besides the photoionization rate, one needs the information of the photon mean free path in order to estimate the reionization source emissivities. At relatively lower redshifts this can be measured by identifying the absorption lines corresponding to the neutral hydrogen structures (Prochaska et al., 2014). However, this identification becomes difficult towards higher redshifts which demands the stacking of the quasar spectra. Recently, relatively short values for the mean free path at $z\sim 6$ have been reported by utilizing this stacking method (Becker et al., 2021; Zhu et al., 2023) which again suggests the scenario of late reionization or excess ionization background fluctuations.
@@
From the discussion so far, it is apparent that we need a reasonably accurate and efficient model for estimating the Ly-$\alpha$ opacity fluctuations during later stages of reionization to properly interpret the observations. However, the modelling faces the challenge of high dynamic range requirements for the simulation box. On the one hand, it has to resolve small scales ($\lesssim 100\leavevmode\nobreak\ \mathrm{Kpc}$) for a converged description of Ly-$\alpha$ forest spectra, while on the other hand, it should have large enough scale to access the observed fluctuation spanning a wide cosmological scale. For example, Davies et al. (2024b) utilizes a very high resolution hydrodynamic simulation, sufficient to resolve small scale structures while their large scale fields are not fully self consistent. In another similar study by Gaikwad et al. (2023), an efficient algorithm has been exploited with a relatively larger simulation box ($160\leavevmode\nobreak\ h^{-1}\mathrm{cMpc}$) and coarser resolution. In principle, there may exist fluctuation features having a length scale larger than these.
@@
In this work, we present an efficient semi numerical model for estimating the large scale Ly-$\alpha$ opacity fluctuations at late stages of reionization. The motivation behind this model is to provide a reasonably accurate description of forest fluctuations within a large cosmological simulation volume while keeping the small scale information intact. We take into account these small scale contributions by properly calibrating the model with respect to a full hydrodynamical simulation. As we shall discuss, the model incorporates other relevant ingredients for large scale opacity fluctuations in a self consistent way. Further, the model also allows us to provide constraints on background mean photoionization rate and an effective mean free path after comparing with the recent Ly-$\alpha$ opacity measurements (Bosman et al., 2022). The method laid out here is a proof of concept study with the aim of efficient EoR inference in the future.
@@ 2.1 Computing Ly-$\alpha$ optical depth
In a semi numerical set up (for examples, Qin et al., 2024; Choudhury & Chakraborty, 2025), the optical depth for Ly-$\alpha$ absorption is generally obtained by Fluctuating Gunn Peterson Approximation (FGPA; Weinberg et al., 1998). The approximation provides a relation between the Ly-$\alpha$ optical depth ($\tau_{\alpha}$) and the underlying density field ($\Delta$) of the IGM i.e.
$\tau_{\alpha}\propto\Delta^{2}\leavevmode\nobreak\ T^{-0.724}\leavevmode\nobreak\ \Gamma_{\mathrm{HI}}^{-1}\leavevmode\nobreak\ (1+z)^{6}\leavevmode\nobreak\ H^{-1}(z)$ (1)
where $T$ and $\Gamma_{\mathrm{HI}}$ are the temperature and photoionization rate with $H(z)$ being the Hubble parametere at redshift $z$. However, this relation is applicable when we can access reasonably small scales ($\sim 10-20\leavevmode\nobreak\ \mathrm{Kpc}$) structures with very high resolution simulation. To apply this on coarse resolution semi numerical simulations, we need to modify the relation after properly calibrating with respect to the high resolution full hydrodynamical simulation. With the aim of calibration, we assume that for each coarse cell ($i$, with $4h^{-1}\mathrm{cMpc}$ resolution in this study), the Ly-$\alpha$ optical depth, $\tau_{i}$ follows power law relations with different physical quantities at the scale of pixel resolution. These power law indices are then calibrated with the post processing estimates of high resolution hydrodynamical simulation. Hence, the modified optical depth for a cell in a fully ionized state (i.e obeying ionizing equilibrium criterion) can be written as
@@
where $\Delta_{i}$, $T_{i}$ and $\Gamma_{\mathrm{HI},i}$ are the overdensity, temperature and photoionization rate for the $i$-th cell respectively with $k$ is a normalization factor. We utilize the full hydrodynamical simulation, Nyx, to estimate the different indices, as we shall discuss in the next paragraph. On the other side, the optical depth for a neutral region can be assumed to be very large, providing essentially zero transmitted flux. Hence, the effective transmitted flux in each cell is given by
$F_{i}=x_{\mathrm{HII},i}e^{-\tau_{i}}$ (3)
The effective optical depth of the box is then computed as
@@
where $n$ is the number of pixels in a single sightline to be averaged over according to the data and delimited-⟨$\langle$..$\rangle$ … ⟩ represents the averaging over all the sightlines. In this study, the averaging length scale is chosen as $\Delta z=0.1$ to be consistent with Ly-$\alpha$ forest data (Bosman et al., 2022) prepared from extended XQR-30 sample.
Figure 1: The figure shows the 2D density plot of optical depth ($\tau_{\mathrm{hyd},i}$) vs density ($\Delta_{i}$) averaged at the pixel scale ($4\leavevmode\nobreak\ h^{-1}\mathrm{cMpc}$) from NyX hydro simulation suite at redshift $z=5.0$. The slope of the correlation matches with $\alpha=0.56$, which is used to calibrate our semi-analytic model.
Figure 2: Global variation of optical depth ($\Delta\log\tau$) at the resolution scale ($4h^{-1}\mathrm{cMpc}$) for variations in mean photoionization rates ($\Delta\log\langle\Gamma_{\mathrm{HI}}\rangle$, in left panel) and for variations in temperatures ($\Delta\log T$, in right panel).
Figure 3: The comparison of optical depths from hydro simulation ($\tau_{\mathrm{hyd}}$) and calibrated empirical model ($\tau_{\mathrm{model}}$) for three different redshifts ($z=5.0,5.5$ $6.0$) used for calibration. The different colors indicate the set of different values of mean photoionization rates.
@@ 2.2 Calibration with Nyx simulation suite
[…]
2.3 Semi numerical setup
In this subsection, we briefly describe semi numerical model setup for computing the different ingredient fields i.e. density, temperature, ultraviolet background (UVB) / photoionization, for estimating the effective optical depths which can be further compared against the observational data.

### v2

# Efficient Modelling of Lyman-$\alpha$ opacity fluctuations during late reionization epoch

@@
Efficient Modelling of Lyman-$\alpha$ opacity fluctuations during late reionization epoch
Barun Maity, maity@mpia.de    Frederick Davies       Prakash Gaikwad
@@
The Lyman-$\alpha$ forest opacity fluctuations observed from high-redshift quasar spectra have been proven to be extremely successful in order to probe the late phase of the reionization epoch. For ideal modeling of these opacity fluctuations, one of the main challenges is to satisfy the extremely high dynamic range requirements of the simulation box, resolving the Lyman-$\alpha$ forest while probing the large cosmological scales. In this study, we adopt an efficient approach to model Lyman-$\alpha$ opacity fluctuations in a coarse simulation volume, utilizing the semi-numerical reionization model SCRIPT (including inhomogeneous recombination and radiative feedback) integrated with a realistic photoionization background fluctuation generating model. Our model crucially incorporates ionization and temperature fluctuations, which are consistent with the reionization model. After calibrating our method with respect to high-resolution full hydrodynamic simulation, Nyx, we compared the models with available observational data at the redshift range, z=5.06.1z=5.0-6.1. With a fiducial reionization model (reionization end at $z=5.8$), we demonstrated that the observed scatter in the effective optical depth can be matched reasonably well by tuning the free parameters of our model, i.e., the effective ionizing photon mean free path and mean photoionization rate. We further pursued an MCMC-based parameter space exploration, utilizing the available data to put constraints on the above free parameters. Our estimation prefers a slightly higher photoionization rate and slightly lower mean free path than the previous studies, which is also a consequence of temperature fluctuations. This study holds significant promise for efficiently extracting important physical information about the Epoch of Reionization, utilizing the wealth of available and upcoming observational data.
@@ 1 Introduction
The Epoch of Reionization (EoR) represents an important phase in the history of the Universe, marking the time when the first stars and galaxies form and ionize neutral hydrogen atoms (HI) in the intergalactic medium (IGM) (see reviews for details, Barkana & Loeb, 2001; Choudhury, 2009; McQuinn, 2016; Dayal & Ferrara, 2018; Gnedin & Madau, 2022; Choudhury, 2022). The precise timeline and morphology of this epoch can be crucial to understand the structure formation and evolution of the universe at those early times. There exists a handful of observational probes at various wavelengths to trace the IGM during the EoR (Zaroubi, 2013; Mesinger, 2016). Among these, spectra obtained from high-redshift quasars have been proven to be a treasure trove in extracting critical information about the nature of the IGM.
@@
Specifically, Lyman-$\alpha$ (Ly-$\alpha$) forest absorption of quasar spectra contains information about the line-of-sight distribution of neutral hydrogen atoms (HI) in the IGM, which has been widely used in the literature to study various physical aspects. For example, these include exploiting realistic numerical simulations to constrain the photoionization rate (Bolton & Haehnelt, 2007; Wyithe & Bolton, 2011; Gaikwad et al., 2023; Davies et al., 2024b) and tracking the temperature evolution of the IGM (Walther et al., 2019; Gaikwad et al., 2020). The high quality of existing quasar spectra also allows us to directly estimate the neutral fraction of the universe with damping wing analysis (Spina et al., 2024) and with dark pixel analysis (McGreer et al., 2015; Jin et al., 2023). Recently, a popular scenario of late reionization end ($z_{\mathrm{end}}\sim 5.3$) has come out based on the findings of large Gunn-Peterson (GP) troughs, giving rise to excess Ly-$\alpha$ forest opacity fluctuations at $z\geq 5.3$ (Zhu et al., 2021, 2022; Bosman et al., 2022).
@@
Several traditional models have been found to be inconsistent in order to explain the large GP trough and the excess Ly-$\alpha$ opacity fluctuations. There exist a few suggestions to match the data, keeping the reionization end at $z\gtrsim 6$. One of these includes a hidden abundance of faint quasar populations at $z\geq 5$ (Meiksin, 2020); however, it is not yet completely understood whether they have the required photon budget to make a significant contribution. An extended and hot reionization scenario can also produce significant fluctuations via the temperature dependence of recombination rate (D’Aloisio et al., 2015). But, this model associates the opaque trough with large-scale galaxy overdensity (Davies et al., 2018a), which is disfavoured by the galaxy survey results (Becker et al., 2018; Christenson et al., 2021). It is also possible that the excess ionizing background fluctuations are amplified due to very short mean free path of the ionizing photons in the IGM (Davies & Furlanetto, 2016; D’Aloisio et al., 2018). However, that doesn’t agree with the extrapolation from existing lower redshift measurements (Worseck et al., 2014). Another emerging explanation involves a late reionization end ($z_{\mathrm{end}}\leq 5.5$) where the shadowing effect on the large-scale ionization fluctuations by the residual neutral islands is responsible for the excess opacity fluctuations (Kulkarni et al., 2019; Keating et al., 2020; Nasir & D’Aloisio, 2020). However, these models are finely tuned to match the ionizing photon emissivity according to the demand of both Ly-$\alpha$ forest fluctuations and other high-redshift observables like CMB scattering optical depth. In recent times, this explanation has also been adopted by semi-numerical models, which are further able to provide constraints on the global ionization fraction utilizing Ly-$\alpha$ forest observations (Choudhury et al., 2021; Qin et al., 2021, 2025). However, these models lack a realistic description of photoionization background fluctuations and temperature fluctuations. Despite recent improvements along these lines (Choudhury & Chakraborty, 2025), those models are still devoid of proper calibration for individual components of opacity fluctuations against a realistic forest model. For instance, Qin et al. (2025) used a single simulation suite at a fixed redshift ($z=5.0$) to map out an analytic relation between true opacity and FGPA approximation via probability density function. However, as we’ll discuss in our study, the FGPA relation itself may need to be modified while dealing with coarse resolution simulation, accounting for the calibration of temperature, UVB fluctuations, and redshift dependence of Ly-$\alpha$ optical depth.
@@
Nevertheless, it is crucial to determine the strength of photoionization background fluctuations as it provides an important boundary limit to the reionization scenario. It has been shown that the fluctuations rise rapidly when the ionization bubbles start to overlap with each other, eliminating the neutral hydrogen patches (Bolton & Haehnelt, 2007; Wyithe & Bolton, 2011; Davies et al., 2018d). The measurements of the photoionization rate ($\Gamma_{\mathrm{HI}}$) also show an increment in the redshift range $z\sim 5-6$ supporting the above statement. However, the constraints are not very stringent due to a lack of transmission flux at these high redshifts and a degeneracy with the thermal state of the IGM. In recent studies, the constraints have become more stringent with the abundance of high-quality data and robust modelling (Gaikwad et al., 2023; Davies et al., 2024b).
@@
Besides the photoionization rate, one needs the information on the photon mean free path in order to estimate the reionization source emissivities. At relatively lower redshifts, this can be measured by identifying the absorption lines corresponding to the neutral hydrogen structures (Prochaska et al., 2014). However, this identification becomes difficult towards higher redshifts, which demands the stacking of the quasar spectra. Recently, relatively short values for the mean free path at $z\sim 6$ have been reported by utilizing this stacking method (Becker et al., 2021; Zhu et al., 2023), which again suggests the scenario of late reionization or excess ionization background fluctuations.
@@
From the discussion so far, it is apparent that we need a reasonably accurate and efficient model for estimating the Ly-$\alpha$ opacity fluctuations during later stages of reionization to properly interpret the observations. However, the modelling faces the challenge of high dynamic range requirements for the simulation box. On the one hand, it has to resolve small scales ($\lesssim 100\penalty 10000\ \mathrm{Kpc}$) for a converged description of Ly-$\alpha$ forest spectra, while on the other hand, it should have a large enough scale to access the observed fluctuation spanning a wide cosmological scale. For example, Davies et al. (2024b) utilizes a very high resolution hydrodynamic simulation, sufficient to resolve small-scale structures while their large-scale fields are not fully self-consistent. In another similar study by Gaikwad et al. (2023), an efficient algorithm has been exploited with a relatively larger simulation box ($160\penalty 10000\ h^{-1}\mathrm{cMpc}$) and coarser resolution. In principle, there may exist fluctuation features having a length scale larger than these.
@@
In this work, we presented an efficient semi-numerical model for estimating the large-scale Ly-$\alpha$ opacity fluctuations at late stages of reionization. The motivation behind this model is to provide a reasonably accurate description of forest fluctuations within a large cosmological simulation volume while keeping the small-scale information intact. We took into account these small-scale contributions by properly calibrating the model with respect to a full hydrodynamical simulation. As we shall discuss, the model incorporated other relevant ingredients for large-scale opacity fluctuations in a self-consistent way. Further, the model also allowed us to provide constraints on background mean photoionization rate and an effective mean free path after comparing with the recent Ly-$\alpha$ opacity measurements (Bosman et al., 2022). The method laid out here is a proof-of-concept study with the aim of efficient EoR inference in the future.
@@ 2.1 Computing Ly-$\alpha$ optical depth
In a semi-numerical set up (for examples, Qin et al., 2025; Choudhury & Chakraborty, 2025), the optical depth for Ly-$\alpha$ absorption is generally obtained by Fluctuating Gunn Peterson Approximation (FGPA; Weinberg et al., 1998). The approximation provides a relation between the Ly-$\alpha$ optical depth ($\tau_{\alpha}$) and the underlying density field ($\Delta$) of the IGM i.e.
$\tau_{\alpha}\propto\Delta^{2}\penalty 10000\ T^{-0.724}\penalty 10000\ \Gamma_{\mathrm{HI}}^{-1}\penalty 10000\ (1+z)^{6}\penalty 10000\ H^{-1}(z)$ (1)
where $T$ and $\Gamma_{\mathrm{HI}}$ are the temperature and photoionization rate with $H(z)$ being the Hubble parameter at redshift $z$. However, this relation is applicable when we can access reasonably small scales ($\sim 10-20\penalty 10000\ \mathrm{Kpc}$) structures with a very high resolution simulation. To apply this on coarse resolution semi-numerical simulations, we need to modify the relation after properly calibrating with respect to the high resolution full hydrodynamical simulation. With the aim of calibration, we assumed that for each coarse cell ($i$, with $4h^{-1}\mathrm{cMpc}$ resolution in this study), the Ly-$\alpha$ optical depth, $\tau_{i}$, follows power law relations with different physical quantities at the scale of pixel resolution. These power law indices are then calibrated with the post-processing estimates of high-resolution hydrodynamical simulations. Hence, the modified optical depth for a cell in a fully ionized state (i.e, obeying ionizing equilibrium criterion) can be written as
@@
where $\Delta_{i}$, $T_{i}$ and $\Gamma_{\mathrm{HI},i}$ are the overdensity, temperature and photoionization rate for the $i$-th cell respectively with $k$ is a normalization factor. We utilized the full hydrodynamical simulation, Nyx, to estimate the different indices, as we shall discuss in the next paragraph. On the other side, the optical depth for a neutral region can be assumed to be very large, providing essentially zero transmitted flux. Hence, the effective transmitted flux in each cell is given by
$F_{i}=x_{\mathrm{HII},i}e^{-\tau_{i}}+x_{\mathrm{HI},i}e^{-\tau_{i}^{n}}\simeq x_{\mathrm{HII},i}e^{-\tau_{i}}$ (3)
where $\tau_{i}^{n}$ is the optical depth for the neutral region, which is sufficiently large and can be approximated to $\infty$. The ionization fraction of $i$-th cell is denoted as $x_{\mathrm{HII},i}=1-x_{\mathrm{HI},i}$. The effective optical depth of the box was then computed as
@@
where $n$ is the number of pixels in a single sightline to be averaged over according to the data, and $\langle$$\rangle$ represents the averaging over all the sightlines. In this study, the averaging length scale was chosen as $\Delta z=0.1$ to be consistent with Ly-$\alpha$ forest data (Bosman et al., 2022) prepared from the extended XQR-30 sample.
Figure 1: The figure shows the 2D density plot of optical depth ($\tau_{\mathrm{hyd},i}$) vs density ($\Delta_{i}$) averaged at the pixel scale ($4\penalty 10000\ h^{-1}\mathrm{cMpc}$) from NyX hydro simulation suite at redshift $z=5.0$. The slope of the correlation matches with $\alpha=0.56$, which is used to calibrate our semi-analytic model.
Figure 2: Global variation of optical depth ($\Delta\log\tau$) at the resolution scale ($4h^{-1}\mathrm{cMpc}$) for variations in mean photoionization rates ($\Delta\log\langle\Gamma_{\mathrm{HI}}\rangle$, in left panel) and for variations in temperatures ($\Delta\log\langle T\rangle$, in right panel).
Figure 3: The comparison of optical depths from hydro simulation ($\tau_{\mathrm{hyd}}$) and calibrated empirical model ($\tau_{\mathrm{model}}$) for three different redshifts ($z=5.0,5.5$ & $6.0$) used for calibration. The different colors indicate the set of different values of mean photoionization rates.
@@ 2.2 Calibration with Nyx simulation suite
As mentioned earlier, we utilized a cosmological hydrodynamical simulation, Nyx, to extract the skewers and calibrate our technique. The simulation description closely follows the previous work by Davies et al. (2024b), which provides sufficient box size ($100\penalty 10000\ h^{-1}\mathrm{Mpc}$) and resolution ($4096^{3}$ dark matter and baryon grids) for converged Ly-$\alpha$ forest statistics. Further, the Ly-$\alpha$ forest opacities are deposited in redshift space using an approximate form for the Voigt profile (Tepper-García, 2006) in 2 km/s grid resolution, given the neutral hydrogen field estimated assuming ionization equilibrium. We used the skewers from this simulation suite at three redshifts ($z=5.0,5.5$ & $6.0$) to calibrate the redshift dependence of our empirical model. For each of these redshifts, we extracted 10,000 random skewers of overdensity and temperature fields from the simulation box. Similarly, we utilized different temperature variants of the simulation suites. The optical depth skewers also include peculiar velocity information, although the effect is supposed to be small, and we did not explicitly take into account that effect in the semi-numerical setup described later. However, due to the large-scale coherence of the velocity field, this effect can be worthwhile to explore in a future study. As this simulation assumes instantaneous reionization, these temperature variations can be obtained by changing the reionization end. We utilized the models with sudden reionization end at $z=6.2,\penalty 10000\ 6.7,\penalty 10000\ 7.7$ & $15$; providing a range of temperature variations where earlier reionization end corresponds to lower temperature and vice versa. We also created a few different variants of effective mean photoionization rate as well by scaling the optical depth skewers in post-processing, where the opacity is inversely related to the photoionization rate.
@@
2.3 Semi-numerical setup
In this subsection, we briefly describe semi-numerical model setup for computing the different ingredient fields, i.e., density, temperature, ultraviolet background (UVB) / photoionization, for estimating the effective optical depths, which can be further compared against the observational data.